# Diff-IP2D: Diffusion-Based Hand-Object Interaction Prediction on Egocentric Videos

### v3

@@ Abstract
Understanding how humans would behave during hand-object interaction (HOI) is vital for applications in service robot manipulation and extended reality. To achieve this, some recent works simultaneously forecast hand trajectories and object affordances on human egocentric videos. The joint prediction serves as a comprehensive representation of future HOI in 2D space, indicating potential human motion and motivation. However, the existing approaches mostly adopt the autoregressive paradigm, which lacks bidirectional constraints within the holistic future sequence, and accumulates errors along the time axis. Meanwhile, they overlook the effect of camera egomotion on first-person view predictions. To address these limitations, we propose a novel diffusion-based HOI prediction method, namely Diff-IP2D, to forecast future hand trajectories and object affordances with bidirectional constraints in an iterative non-autoregressive manner on egocentric videos. Motion features are further integrated into the conditional denoising process to enable Diff-IP2D aware of the camera wearer’s dynamics for more accurate interaction prediction. Extensive experiments demonstrate that Diff-IP2D significantly outperforms the state-of-the-art baselines on both the off-the-shelf and our newly proposed evaluation metrics. This highlights the efficacy of leveraging a generative paradigm for 2D HOI prediction. The code of Diff-IP2D will be released as open source at https://github.com/IRMVLab/Diff-IP2D.
@@ I Introduction
Accurately anticipating human intentions and future actions is important for artificial intelligence systems in robotics and extended reality [1, 2, 3]. Recent works have tried to tackle the problem from various perspectives, including action recognition and anticipation [4, 5, 6, 7], gaze prediction [8, 9, 10, 11], hand trajectory prediction [12, 13, 14, 15], and object affordance extraction [12, 16, 14, 17]. Among them, jointly predicting hand motion and object affordances can effectively facilitate more reasonable robot manipulation as the prior contextual information, which has been demonstrated on some robot platforms [1, 18, 19]. We believe that deploying such models pretrained by internet-scale human videos on robots is a promising path towards embodied agents. Therefore, our work aims to jointly predict hand trajectories and object affordances on egocentric videos as a concrete hand-object interaction (HOI) expression, following the problem modeling of previous works [12, 14].
@@
Figure 1: Diff-IP2D vs. Existing Paradigm. The existing HOI prediction paradigm (a) tends to accumulate prediction errors under unidirectional constraints. In contrast, our proposed Diff-IP2D (b) directly forecasts all future HOI states in parallel with denoising diffusion, mitigating error accumulation with bidirectional constraints.
@@
Currently, the state-of-the-art approaches [12, 13] predicting hand trajectories and object affordances on egocentric videos tend to exploit the autoregressive (AR) model. They reason about the next HOI state only according to the previous steps (Fig. 1(a)). However, expected “post-contact states” also affect “pre-contact states” according to human intentions that persist across the holistic HOI process as an oracle. Inspired by this, we argue that predicting future HOI states in parallel considering the bidirectional constraints within the holistic sequence outperforms generating the next state autoregressively. With diffusion models emerging across multiple domains [20, 21, 22, 23, 24, 25, 26, 27], their strong forecasting capability has been widely validated. Therefore, we propose a diffusion-based method to predict future hand-object interaction in parallel, considering bidirectional constraints in the latent space compared to the traditional autoregressive generation (Fig. 1(b)).
@@
Moreover, we also identify two inherent gaps affecting HOI prediction in the existing paradigm: 1) Directly predicting the projection of 3D future hand trajectories and object affordances on 2D egocentric image plane is an ill-posed problem involving spatial ambiguities. There is generally a gap between 2D pixel movements and 3D real actions, which can be bridged by spatial transformation across multiple views changing with egomotion. 2) The past egocentric videos are absorbed to predict future interaction states on the last observed image, which is actually a “canvas” from a different view w.r.t all the other frames. Therefore, there is also a gap between the last observation (egocentric view) and the other observations (analogous to exocentric view) caused by egomotion. To fill the two gaps together, we further propose to integrate the camera wearer’s egomotion into our diffusion-based paradigm. It enables the denoising model aware of the camera wearer’s dynamics and the spatial relationship between consecutive frames.
@@
The main contributions of this paper are as follows: 1) We propose a diffusion-based hand-object interaction prediction method, dubbed Diff-IP2D. To our best knowledge, this is the first work using the devised denoising diffusion probabilistic model to jointly forecast future hand trajectories and object affordances with only 2D egocentric videos as input. It provides a foundation generative paradigm in the field of HOI prediction. 2) The homography egomotion features are integrated to fill the motion-related gaps inherent in HOI prediction on egocentric videos. 3) We extend the existing metrics and propose the first protocol for jointly evaluating the performance of hand trajectory prediction and object affordance prediction. Comprehensive experiments are conducted to demonstrate that our Diff-IP2D can predict better hand trajectories and object affordances compared to the state-of-the-art baselines, showing its potential for deployment on artificial intelligence systems.
@@ II Related work
Predicting hand-object interaction. Analyzing only past human behavior may be insufficient for service robot manipulation or extended reality. Forecasting possible future object-centric HOI states based on historical observations is also valuable, which attracts increasing attention due to the general knowledge that can be transferred to robot applications [1, 18, 19, 44]. For example, Dessalene et al. [45] propose to generate contact anticipation maps and next active object segmentations as future HOI predictions. Liu et al. [14] first achieve hand trajectory and object affordance prediction simultaneously, revealing that predicting hand motion benefits the extraction of interaction hotspots, and vice versa. Following this work, Liu et al. [12] further develop an object-centric Transformer to jointly forecast future trajectories and affordances autoregressively, and annotate publicly available datasets to support future works. More recently, Bao et al. [13] lift the problem to 3D spaces where hand trajectories are predicted by an uncertainty-aware state space Transformer in an autoregressive manner. However, it needs additional 3D perception inputs from the RGB-D camera. In this work, we still achieve joint hand trajectory and object affordance prediction on 2D human videos rather than in 3D space. We focus on capturing more general knowledge from only egocentric 2D observations in an iterative non-autoregressive (iter-NAR) manner, rather than the autoregressive way of the state-of-the-art works [12, 13].
@@
Figure 2: System Overview of Diff-IP2D. Our proposed paradigm takes in sequential past egocentric images and jointly predicts hand trajectories and object affordances as future HOI states. The observations are mapped to the latent feature space for the diffusion process.
@@ III-A Preliminaries
Task definition. Given the video clip of past egocentric observations =$\mathcal{I}=\{I_{t}\}_{t=-N_{\text{p}}+1}^{0}$, we aim to predict future hand trajectories =$\mathcal{H}=\{H_{t}^{\scriptstyle{\text{R}}},H_{t}^{\scriptstyle{\text{L}}}\}_{t=1}^{N_{\text{f}}}(H_{t}^{\scriptstyle{\text{R}}},H_{t}^{\scriptstyle{\text{L}}}\in\mathbb{R}^{2})$ and potential object contact points $\mathcal{O}=\{O_{n}\}_{n=1}^{N_{\text{o}}}(O_{n}\in\mathbb{R}^{2})$, where $N_{\text{p}}$ and $N_{\text{f}}$ are the numbers of frames in the past and future time horizons respectively, and $N_{\text{o}}$ denotes the number of predicted contact points used to calculate interaction hotspots as object affordances. Following the previous works [12, 14], we predict the future positions of the right hand, the left hand, and the affordance of the next active object on the last observed image of the input videos as a canvas.
Diffusion models. In this work, we propose a diffusion-based approach to gradually corrupt the input to noisy features and then train a denoising model to reverse this process. We first map the input images into a latent space $\mathbf{z}_{0}\sim q(\mathbf{z}_{0})$, which is then corrupted to a standard Gaussian noise $\mathbf{z}_{S}\sim\mathcal{N}(0,\textbf{\text{I}})$. In the forward process, the perturbation operation can be represented as $q(\mathbf{z}_{s}|\mathbf{z}_{s-1})=\mathcal{N}(\mathbf{z}_{s};\sqrt{1-\beta_{s}}\mathbf{z}_{s-1},\beta_{s}\textbf{\text{I}})$, where $\beta$ is the predefined variance scales. In the reverse process, we set a denoising diffusion model to gradually reconstruct the latent $\mathbf{z}_{0}$ from the noisy $\mathbf{z}_{S}$. The denoised features can be used to recover the final future hand trajectories and object affordances.
@@ III-B Architecture
System overview. Accurately reconstructing the future part of the input sequence is critical in diffusion-based prediction. We found that ground-truth hand waypoints $\mathcal{H}^{\text{gt}}=\{H_{t}^{{\scriptstyle{\text{R}}},\text{gt}},H_{t}^{{\scriptstyle{\text{L}}},\text{gt}}\}_{t=1}^{N_{\text{f}}}(H_{t}^{{\scriptstyle{\text{R}}},\text{gt}},H_{t}^{{\scriptstyle{\text{L}}},\text{gt}}\in\mathbb{R}^{2})$ and contact points $\mathcal{O}^{\text{gt}}=\{O^{\text{gt}}_{n}\}_{n=1}^{N_{\text{o}}}(O^{\text{gt}}_{n}\in\mathbb{R}^{2})$ provide discrete and sparse supervision for reconstruction, which is not enough for capturing high-level semantics such as human intentions in the denoising process. Therefore, as Fig. 2 shows, we first use the Multi-Feature Extractor and Side-Oriented Fusion Module to transform input images into latent HOI features, and then implement diffusion-related operation in the latent continuous space. The HOI features denoised by the Motion-Aware Denoising Transformer are further absorbed by the Hand Trajectory Head and Object Affordance Head to generate future hand trajectories and object hotspots.
@@
Multi-Feature Extractor (MFE). Following the previous work [12], we use MFE that consists of a pretrained Temporal Segment Network (TSN) provided by Furnari et al. [34], RoIAlign [52] with average pooling, and Multilayer Perceptron (MLP) to extract hand, object, and global intermediate features for each sequence image $I_{t}\in\mathcal{I}$. The positions of hand-object bounding boxes detected by the off-the-shelf approach [15] are also encoded to feature vectors fused with hand and object intermediate features.
@@
Side-Oriented Fusion Module (SOFM). Our proposed SOFM is a learnable linear transformation to fuse the above-mentioned three types of feature vectors into the final latent form for two sides respectively. Specifically, the global features and right-side features (right-hand/object features) are concatenated to the right-side HOI features $\mathcal{F}^{\scriptstyle{\text{R}}}=\{F_{t}^{\scriptstyle{\text{R}}}\}_{t=-N_{\text{p}}+1}^{X}(F_{t}^{\scriptstyle{\text{R}}}\in\mathbb{R}^{a},X=N_{\text{f}}\,\,\text{for training and}\,\,X=0\,\,\text{for inference})$. The operation and feature sizes are the same as the left-side counterparts, leading to $\mathcal{F}^{\scriptstyle{\text{L}}}=\{F_{t}^{\scriptstyle{\text{L}}}\}_{t=-N_{\text{p}}+1}^{X}$. We further concatenate the side-oriented features along the time axis respectively to generate the input latents $F^{\scriptstyle{\text{R}}}_{\text{seq}}$, $F^{\scriptstyle{\text{L}}}_{\text{seq}}\in\mathbb{R}^{(N_{\text{p}}+X)\times a}$ for the following diffusion model.
@@
Figure 3: Architecture of our proposed MADT. MADT receives corrupted latent HOI features with the position embedding (PE) and time embedding (TE), and outputs denoised future HOI features under the egomotion guidance.
@@
Motion-Aware Denoising Transformer (MADT). Our proposed MADT takes in the noisy latent HOI features and reconstructs future HOI features for the following predictors conditioned on past HOI counterparts. MADT consists of devised Transformer layers as shown in Fig. 3, thus imposing bidirectional constraints on temporal latents. Following the previous work [26], we anchor the past HOI features for both forward and reverse processes. We only impose noises and denoise at the positions of the future feature sequence. The features of the two sides are denoised using the same model, leading to $\hat{F}^{\scriptstyle{\text{R}}}_{\text{seq}}$ and $\hat{F}^{\scriptstyle{\text{L}}}_{\text{seq}}$. In addition, egomotion guidance is proposed here to fill the gaps mentioned in Sec. I. Specifically, we first extract SIFT [53] descriptors to find the pixel correspondence between two adjacent past images in $\mathcal{I}$. Then we use RANSAC [54] to solve the homography matrix that maximizes the number of inliers in keypoint pairs. We accumulate the consecutive homography matrices and obtain $M_{\text{seq}}\in\mathbb{R}^{N_{\text{p}}\times 3\times 3}$ representing the camera wearer’s motion between $I_{t}$ ($t\leq 0$) and $I_{0}$. They are further linearly embedded into an egomotion feature $E_{\text{seq}}\in\mathbb{R}^{N_{\text{p}}\times b}$ by Motion Encoder. The multi-head cross-attention module (MHCA) in MADT then absorbs the egomotion features to guide the denoising process.
@@
Predictors. Our proposed predictors consist of the Hand Trajectory Head (HTH) and Object Affordance Head (OAH). HTH contains an MLP that receives the future parts of the denoised features, $\hat{F}^{\scriptstyle{\text{R}}}_{\text{seq}}[N_{\text{p}}\!+\!1\!:\!N_{\text{p}}\!+\!N_{\text{f}}]$ and $\hat{F}^{\scriptstyle{\text{L}}}_{\text{seq}}[N_{\text{p}}\!+\!1\!:\!N_{\text{p}}\!+\!N_{\text{f}}]$ to generate future waypoints $\mathcal{H}$ of two hands. As to OAH, we empirically exploit Conditional Variational Autoencoder (C-VAE) [55] to generate possible contact points $\mathcal{O}$ in the near future. Taking the right hand as an example, the condition is selected as the time-averaged $\hat{F}^{\scriptstyle{\text{R}}}_{\text{seq}}$ and predicted waypoints $H_{t}^{\scriptstyle{\text{R}}}$. Note that we additionally consider denoised future HOI features $\hat{F}^{\scriptstyle{\text{R}}}_{\text{seq}}[N_{\text{p}}\!+\!1:N_{\text{p}}\!+\!N_{\text{f}}]$ ($t\!>\!0$) besides the features from the past observation ($t\!\leq\!0$) for object affordance prediction. This aligns with the intuitive relationship between the contact points and the overall interaction process.

### v4

@@ Abstract
Understanding how humans would behave during hand-object interaction (HOI) is vital for applications in service robot manipulation and extended reality. To achieve this, some recent works simultaneously forecast hand trajectories and object affordances on human egocentric videos. The joint prediction serves as a comprehensive representation of future HOI in 2D space, indicating potential human motion and motivation. However, the existing approaches mostly adopt the autoregressive paradigm, which lacks bidirectional constraints within the holistic future sequence, and accumulates errors along the time axis. Meanwhile, they overlook the effect of camera egomotion on first-person view predictions. To address these limitations, we propose a novel diffusion-based HOI prediction method, namely Diff-IP2D, to forecast future hand trajectories and object affordances with bidirectional constraints in an iterative non-autoregressive manner on egocentric videos. Motion features are further integrated into the conditional denoising process to enable Diff-IP2D aware of the camera wearer’s dynamics for more accurate interaction prediction. Extensive experiments demonstrate that Diff-IP2D significantly outperforms the state-of-the-art baselines on both the off-the-shelf and our newly proposed evaluation metrics. This highlights the efficacy of leveraging a generative paradigm for 2D HOI prediction. The code of Diff-IP2D is released as open source at https://github.com/IRMVLab/Diff-IP2D.
@@ I Introduction
Accurately anticipating human intentions and future actions is important for artificial intelligence systems in robotics and extended reality [1, 2, 3]. Recent works have tried to tackle the problem from various perspectives, including action recognition and anticipation [4, 5, 6], gaze prediction [7, 8, 9], hand trajectory prediction [10, 11, 12, 13], and object affordance extraction [10, 14, 12, 15]. Among them, jointly predicting hand motion and object affordances can effectively facilitate more reasonable robot manipulation as the prior contextual information, which has been demonstrated on some robot platforms [1, 16, 17]. We believe that deploying such models pretrained by internet-scale human videos on robots is a promising path towards embodied agents. Therefore, our work aims to jointly predict hand trajectories and object affordances on egocentric videos as a concrete hand-object interaction (HOI) expression, following the problem modeling of the previous works [10, 12].
@@
Figure I: Diff-IP2D vs. Existing Paradigm. The existing HOI prediction paradigm (a) tends to accumulate prediction errors under unidirectional constraints. In contrast, our proposed Diff-IP2D (b) directly forecasts all future HOI states in parallel with motion-aware denoising diffusion, mitigating error accumulation with bidirectional constraints.
@@
Currently, the state-of-the-art (SOTA) approaches [10, 11] predicting hand trajectories and object affordances on egocentric videos tend to exploit the autoregressive (AR) model. They reason about the next HOI state only according to the previous steps (Fig. I(a)), focusing on unidirectionally forward constraints with temporal causality. However, expected “post-contact states” also affect “pre-contact states” following spatial causality with human intentions that persist across the holistic HOI process as an oracle (also refer to the cup example in Sec. 2 of the supplementary material). Inspired by this, we argue that predicting future HOI states in parallel attending to bidirectional constraints within the holistic sequence outperforms generating the next state autoregressively and mitigates temporal error accumulation. With diffusion models emerging across multiple domains [18, 19, 20, 21, 22, 23, 24, 25], their strong forecasting capability has been widely validated. Therefore, we propose a diffusion-based method to predict future hand-object interaction in an iterative non-autoregressive (iter-NAR) manner (Fig. I(b)), considering bidirectional constraints in the latent space compared to traditional AR generation.
@@
Moreover, we also find two inherent gaps affecting HOI prediction in the existing paradigm: 1) Directly predicting the projection of 3D future hand trajectories and object affordances on 2D egocentric image plane is an ill-posed problem involving spatial ambiguities. There is generally a gap between 2D pixel movements and 3D real actions, which can be bridged by spatial transformation across multiple views changing with egomotion. 2) The past egocentric videos are absorbed to predict future interaction states on the last observed image, which is actually a “canvas” from a different view w.r.t all the other frames. Therefore, there is also a gap between the last observation (egocentric view) and the other observations (analogous to exocentric view) caused by egomotion. To fill these two gaps, we further propose integrating the camera wearer’s egomotion into our diffusion-based paradigm by the cross-attention between egomotion homography features and HOI features. It enables the denoising model aware of the camera wearer’s dynamics and the spatial relationship between consecutive frames.
@@
The main contributions of this paper are as follows: 1) We propose a diffusion-based hand-object interaction prediction method, dubbed Diff-IP2D. To our best knowledge, this is the first work using the devised denoising diffusion probabilistic model to jointly forecast future hand trajectories and object affordances with only 2D egocentric videos as input. It provides a foundation iter-NAR generative paradigm in the field of HOI prediction. 2) The homography egomotion features are integrated to fill the motion-related gaps inherent in HOI prediction on egocentric videos. 3) Comprehensive experiments conducted on our extended evaluation metrics demonstrate that Diff-IP2D can predict better hand trajectories and object affordances compared to the SOTA baselines, showing its potential for deployment on artificial intelligence systems.
@@ II Related work
Predicting hand-object interaction. Analyzing only past human behavior may be insufficient for service robot manipulation or extended reality. Forecasting possible future object-centric HOI states based on historical observations is also valuable, which attracts increasing attention due to the general knowledge that can be transferred to robot applications [1, 16, 17, 42]. For example, Dessalene et al. [43] propose to generate contact anticipation maps and next active object segmentations as future HOI predictions. Liu et al. [12] first achieve hand trajectory and object affordance prediction simultaneously, revealing that predicting hand motion benefits the extraction of interaction hotspots, and vice versa. Following this work, Liu et al. [10] further develop an object-centric Transformer to jointly forecast future trajectories and affordances autoregressively, and annotate publicly available datasets to support future works. More recently, Bao et al. [11] lift the problem to 3D spaces where hand trajectories are predicted by an uncertainty-aware state space Transformer in an autoregressive manner. However, it needs additional 3D perception inputs from the RGB-D camera. In this work, we still achieve joint hand trajectory and object affordance prediction on 2D human videos rather than in 3D space. We focus on capturing more general knowledge from only egocentric 2D observations in an iterative non-autoregressive manner, rather than the autoregressive way of the SOTA works [10, 11].
@@
Figure II: System overview of Diff-IP2D. Our proposed paradigm takes in sequential past egocentric images and jointly predicts hand trajectories and object affordances as future HOI states. The observations are mapped to the latent feature space for the diffusion process.
@@ III Proposed Method
In this section, we first introduce the preliminaries in Sec. III-A. Then we elaborate on our proposed Diff-IP2D architecture in Sec. III-B. Next, we clarify the training and inference schemes in Sec. III-C and Sec. III-D respectively.
@@ III-A Preliminaries
Task definition. Given the video clip of past egocentric observations =$\mathcal{I}=\{I_{t}\}_{t=-N_{\text{p}}+1}^{0}$, we aim to predict future hand trajectories =$\mathcal{H}=\{H_{t}^{\scriptstyle{\text{R}}},H_{t}^{\scriptstyle{\text{L}}}\}_{t=1}^{N_{\text{f}}}(H_{t}^{\scriptstyle{\text{R}}},H_{t}^{\scriptstyle{\text{L}}}\in\mathbb{R}^{2})$ and potential object contact points $\mathcal{O}=\{O_{n}\}_{n=1}^{N_{\text{o}}}(O_{n}\in\mathbb{R}^{2})$, where $N_{\text{p}}$ and $N_{\text{f}}$ are the numbers of frames in the past and future time horizons respectively, and $N_{\text{o}}$ denotes the number of predicted contact points used to calculate interaction hotspots as object affordances. Following the previous works [10, 12], we predict future positions of the right hand, left hand, and affordance of the next active object on the last observed image of the input videos as a canvas.
Diffusion models. In this work, we propose a diffusion-based approach to gradually corrupt the input to noisy features and then train a denoising model to reverse this process. We first map the input images into a latent space $\mathbf{z}_{0}\sim q(\mathbf{z}_{0})$, which is then corrupted to a standard Gaussian noise $\mathbf{z}_{S}\sim\mathcal{N}(0,\textbf{\text{I}})$. In the forward process, the perturbation operation can be represented as $q(\mathbf{z}_{s}|\mathbf{z}_{s-1})=\mathcal{N}(\mathbf{z}_{s};\sqrt{1-\beta_{s}}\mathbf{z}_{s-1},\beta_{s}\textbf{\text{I}})$, where $\beta_{s}$ is the predefined variance scale. In the reverse process, we set a denoising diffusion model to gradually reconstruct the latent $\mathbf{z}_{0}$ from the noisy $\mathbf{z}_{S}$. The denoised features can be used to recover the final future hand trajectories and object affordances.
@@ III-B Architecture
System overview. Accurately reconstructing the future part of the input sequence is critical in diffusion-based prediction. We found that ground-truth hand waypoints $\mathcal{H}^{\text{gt}}=\{H_{t}^{{\scriptstyle{\text{R}}},\text{gt}},H_{t}^{{\scriptstyle{\text{L}}},\text{gt}}\}_{t=1}^{N_{\text{f}}}(H_{t}^{{\scriptstyle{\text{R}}},\text{gt}},H_{t}^{{\scriptstyle{\text{L}}},\text{gt}}\in\mathbb{R}^{2})$ and contact points $\mathcal{O}^{\text{gt}}=\{O^{\text{gt}}_{n}\}_{n=1}^{N_{\text{o}}}(O^{\text{gt}}_{n}\in\mathbb{R}^{2})$ provide discrete and sparse supervision for reconstruction, which is not enough for capturing high-level semantics such as human intentions in the denoising process. Therefore, as Fig. II shows, we first use the Multi-Feature Extractor (MFE) and Side-Oriented Fusion Module (SOFM) to transform input images into latent HOI features, and then implement diffusion-related operation in the latent continuous space. The HOI features denoised by the Motion-Aware Denoising Transformer (MADT) are further absorbed by the Hand Trajectory Head and Object Affordance Head to generate future hand trajectories and object hotspots.
@@
Multi-Feature Extractor (MFE). Following the previous work [10], we use MFE that consists of a pretrained Temporal Segment Network (TSN) provided by Furnari et al. [32], RoIAlign [50] with average pooling, and Multilayer Perceptron (MLP) to extract hand, object, and global intermediate features for each sequence image $I_{t}\in\mathcal{I}$. The positions of hand-object bounding boxes detected by the off-the-shelf approach [13] are also encoded to feature vectors fused with hand and object intermediate features. That is, all the following HOI features in this work encompass spatial information of hands and objects within each image.
@@
Side-Oriented Fusion Module (SOFM). Our proposed SOFM is a learnable linear transformation to fuse the above-mentioned three types of feature vectors into the final latent form for two sides respectively. Specifically, the global features and right-side features (right-hand/object features) are concatenated and are then linearly transformed to the right-side HOI features $\mathcal{F}^{\scriptstyle{\text{R}}}=\{F_{t}^{\scriptstyle{\text{R}}}\}_{t=-N_{\text{p}}+1}^{X}(F_{t}^{\scriptstyle{\text{R}}}\in\mathbb{R}^{a},X=N_{\text{f}}$ for training and $X=0$ for inference). The operation and feature sizes are the same as the left-side counterparts, leading to $\mathcal{F}^{\scriptstyle{\text{L}}}=\{F_{t}^{\scriptstyle{\text{L}}}\}_{t=-N_{\text{p}}+1}^{X}$. We further concatenate the side-oriented features along the time axis respectively to generate the input latents $F^{\scriptstyle{\text{R}}}_{\text{seq}}$, $F^{\scriptstyle{\text{L}}}_{\text{seq}}\in\mathbb{R}^{(N_{\text{p}}+X)\times a}$ for the following diffusion model.
@@
Figure III: MADT architecture. It receives corrupted HOI latents with position embedding (PE) and time embedding (TE), and outputs denoised future HOI latents under egomotion guidance.
@@
Motion-Aware Denoising Transformer (MADT). Our proposed MADT takes in the noisy latent HOI features and reconstructs future HOI features for the following predictors conditioned on past HOI counterparts. MADT consists of devised Transformer layers as shown in Fig. III, thus imposing bidirectional constraints on temporal latents. Following the previous work [24], we anchor the past HOI features for both forward and reverse processes. We only impose noises and denoise at the positions of the future feature sequence. The features of the two sides are denoised using the same model, leading to $\hat{F}^{\scriptstyle{\text{R}}}_{\text{seq}}$ and $\hat{F}^{\scriptstyle{\text{L}}}_{\text{seq}}$. In addition, egomotion guidance is proposed here to fill the gaps mentioned in Sec. I. Specifically, we first extract SIFT [51] descriptors to find the pixel correspondence between two adjacent past images in $\mathcal{I}$. Then we use RANSAC [52] to solve the homography matrix that maximizes the number of inliers in keypoint pairs. We accumulate the consecutive homography matrices and obtain $M_{\text{seq}}\in\mathbb{R}^{N_{\text{p}}\times 3\times 3}$ representing the camera wearer’s motion between $I_{t}$ ($t\leq 0$) and $I_{0}$. They are further linearly embedded into an egomotion feature $E_{\text{seq}}\in\mathbb{R}^{N_{\text{p}}\times b}$ by Motion Encoder. The multi-head cross-attention module (MHCA) in MADT then absorbs the egomotion features to guide the denoising process. More analysis on the use of egomotion guidance can be found in Sec. 1 of the supplementary material.
@@
Predictors. Our proposed predictors consist of Hand Trajectory Head (HTH) and Object Affordance Head (OAH). HTH contains an MLP that receives the future parts of the denoised features $\hat{F}^{\scriptstyle{\text{R}}}_{\text{seq}}[N_{\text{p}}\!+\!1\!:\!N_{\text{p}}\!+\!N_{\text{f}}]$ and $\hat{F}^{\scriptstyle{\text{L}}}_{\text{seq}}[N_{\text{p}}\!+\!1\!:\!N_{\text{p}}\!+\!N_{\text{f}}]$, to generate future hand waypoints $\mathcal{H}$ of two hands. As to OAH, we empirically exploit Conditional Variational Autoencoder (C-VAE) [53] to generate possible contact points $\mathcal{O}$ of the next active object in near future. Taking the right hand as an example, the condition is selected as the time-averaged $\hat{F}^{\scriptstyle{\text{R}}}_{\text{seq}}$ and predicted waypoints $H_{t}^{\scriptstyle{\text{R}}}$. Note that we additionally consider denoised future HOI features $\hat{F}^{\scriptstyle{\text{R}}}_{\text{seq}}[N_{\text{p}}\!+\!1:N_{\text{p}}\!+\!N_{\text{f}}]$ ($t\!>\!0$) besides the features from past observations ($t\!\leq\!0$) for object affordance prediction. This aligns with the intuitive relationship between the contact points and the overall interaction process.